OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2? y = ? x + ?

OpenStudy (alexandervonhumboldt2):

slope is the coefficent with x y-intersect is the second value

OpenStudy (alexandervonhumboldt2):

y=-2x+b 12=-16+b b=28 y=-2x+28

OpenStudy (anonymous):

you lost me?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (anonymous):

no i mean i don't understand

OpenStudy (alexandervonhumboldt2):

so fist write y=-2x+b now substitute coordinates of point: 12=-2*8+b solve for b

OpenStudy (alexandervonhumboldt2):

b=28 so wiite y=-2x+28

OpenStudy (anonymous):

nevermind i thought you wrote something else cause i have dyslexia sorry.

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (anonymous):

Thanks!